OpenStudy (anonymous):

find the equation of the tangents through the point (8,-1) to the ellipse 2x^2+5y^2=70

OpenStudy (tkhunny):

1) Check that it is possible. Is (8,-1) actually OUTSIDE the ellipse?

OpenStudy (solomonzelman):

Yes, (8,-1) is not on the ellipse, because \(5y^2\) is odd at y=-1, and \(2x^2\) is even at any integer x (certain at an even x=8). Even + Odd \(\ne\) Even

OpenStudy (anonymous):

my graphing calculator gives this 

OpenStudy (solomonzelman):

So you have two tangents 

OpenStudy (tkhunny):

As in, OUTSIDE, not just "not on".

OpenStudy (tkhunny):

Given some other point, other than (8,-1), say (a,b), what is the equation of the line containing both (8,1) and (a,b)?

OpenStudy (anonymous):

the slope would be \[\frac{ b+1 }{ a-8 }\]so the equation would be y+1=(b+1/a-8)*(x-8)

OpenStudy (anonymous):

like that?

OpenStudy (mathmate):

@aliLnn Let's first put the equation of the ellipse in order: 2x^2+5y^2=70 => \(\large \frac{x^2}{35}+\frac{y^2}{14}=1\)...................(1) => \(a^2=35, b^2=14\)...............(2) Then, you can assume the line to have slope m, so the line passing through (8,-1)=(x0,y0) is y-y0=m(x-x0), or y=mx+(y0-mx0) \(y= mx+q\)..............................(3) where \(q=(y0-mx0)\)..............(4) At the point of tangency, it's a common point between the line and the ellipse, so we can substitute (3) into (1) and solve for x to get \((a^2m^2+b^2)x^2+2a^2mqx+a^2q^2-a^2b^2=0\)............(5) Let \(A=a^2m^2+b^2\) ..........(6a) \(B=2a^2mq\), and..............(6b) \(C=a^2q^2-a^2b^2\)...............(6c) (5) simplifies to \(Ax^2+Bx+C=0\)...............................................................(5a) The condition for tangency is such that there is a double root (multiplicity 2), or B^2-4AC=0 ......................................(7) Substitute (6a), (6b) and (6c) into (7) results in \(-4a^2b^2(q^2-a^2m^2-b^2)=0\) ...........................(8) Since \(a^2\ne 0\) and \(b^2\ne 0\), we have \(q^2-a^2m^2-b^2=0\) ...........................................(8a) from which we can solve for m (a^2, b^2, x0, y0 are all numeric quantities). Can you solve the quadratic (8a) for m? There are two distinct roots.

OpenStudy (mathmate):

By the way, q=q(m), so substitute equation (4) in (8a) before solving for m. For checking: the solutions for m satisfy |m|\(\le\)1, or you can post your answers for checking. After that, you would use equation (3) to find the tangent lines.